OpenStudy (anonymous):

Suppose that y varies inversely with x, and y = 6 when x = 8. What is an equation for the inverse variation? y = x/48 y = 2x x = y/2 y = 48/x

OpenStudy (anonymous):

Help! (:

OpenStudy (anonymous):

Well, The statement " y varies inversely as x means that when x increases, y decreases by the same factor. In other words, the expression xy is constant: xy = k where k is the constant of variation. So, the relationship between x and y as: y=k/x Do with the formula :)

OpenStudy (anonymous):

@Machida this sort of thing really confuses me.. what would the answer be? :)

OpenStudy (anonymous):

The last one since y=1/x

OpenStudy (anonymous):

@Thefaceless thanks!! :)

OpenStudy (anonymous):

y = 6 x = 8 then k=xy find the k first. :)

OpenStudy (anonymous):

@Machida is it correct it would be the last one??(:

OpenStudy (anonymous):

when use y=k/x , it will be y=48/x

OpenStudy (anonymous):

Yes, because y=48/x ---> xy=48 ----> (6)(8)=48

OpenStudy (anonymous):

Nah :)

OpenStudy (anonymous):

@Machida are you sure??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Im sure :)

OpenStudy (anonymous):

@Machida so its y=48/x??

OpenStudy (anonymous):

lol yes.

OpenStudy (anonymous):

Yes. smh.

OpenStudy (anonymous):

@machida (x thanks so mcuh!